I need help, PLEASE~~
t/3-t+7<-4t/3+8

that is supposed to be is less then or equal, too, i think i didnt know how to put the line under it!

Answer by scott8148(5882)

combining the t terms on the left gives -2t/3+7<=-4t/3+8 ... subtracting 7 gives -2t/3<=-4t/3+1

multiplying by 3 gives -2t<=-4t+3 ... adding 4t gives 2t<=3 ... dividing by 2 gives t<=3/2